the incidence of rheumatoid arthritis is about 4/100,000 per year and the prevalence about 1/1000 persons on average how many years does the disease persist
Answer #1

Prevalence= 1/1000 persons

Incidence = 4/100,000 per year

Prevalence= (Incidence Rate) X (Average Duration of disease)

Duration of disease= Prevalence/ Incidence rate

Duration of disease = (1/1000) / (4/100,000)

=0.001 /0.00004

Duration of disease =25 years.